New Tab SMOLIRA GOLF CORP 2014 and 2015 Liabilities and Owners Equity 2014 2014 2015 2015 Current assets Current liabilities Cash Accounts receivable Inventory s 24,08624,500 15,600 27,500 12,848 26,042 Notes payable Other S 23,584 15,000 11,971 27,500 11,200 18,300 Total s 62.976 67,600 Total $ 50,555 57,000 S 60,000 Owners equity Common stock and paid-in surplus 44,000 44,000 256,000 237,116 Net plant and equipment $328 695 362400 Total $391,671 $430.000 Total S281,116 $300,000 5391671 430.000 Total assets SMOLIRA GOLF CORP 2015 Income Statement Sales Cost of goods sold Depreciation 254,500 32,800 s 86.173 14,700 interest paid 5 42 884

Smolira Golf Corp. has 40,000 shares of common stock outstanding, and the market price for a share of stock at the end of 2015 was $30.

What is the market to book ratio at the end of 2015?
Answer #1
Book value per share = Total Equity/Shares outstanding = 300000/40000 = $7.5
Market to book ratio at the end of 2015 = Market price/Book value = 30/7.5 = 4